OpenStudy (anonymous):

find solutions to the system of equations: y=x^2-5x+3 and y= -x-2

OpenStudy (pottersheep):

You have to equate them to each other

OpenStudy (pottersheep):

so x^2 -5x+3 = -x - 2 then you solve

OpenStudy (pottersheep):

Ok?

OpenStudy (anonymous):

once I have x^2=4x-5 what do I do